The diameter of a circle whose area is approximately 113.04cm squared
Answer by mananth(16946)

area = 113.04
pi*r^2= 113.04
r^2=113.04/pi
r = sqrt ( 113.04/pi)
r=6
so diameter = 12 cm
...
m.ananth@hotmail.ca